# CASteer: Cross-Attention Steering for Controllable Concept Erasure

### v3

@@ F.1 Experimental setup
F.2 Quantitative results
In this section, we present quantitative results on steering SANA model. We use the same experimental setups as for SD-1.4, described in Sec.4. Results on SANA model are shown in Tab.12,13,14,15.
Table 12: Quantitative results on nudity removal based on I2P DBLP:conf/cvpr/SchramowskiBDK23 dataset. Detection of nude body parts is done by Nudenet at a threshold of 0.6. F: Female, M: Male. The best results are highlighted in bold, second-best are underlined.
@@ F.2 Quantitative results
Table 13: Quantitative results on inappropriate content removal based on I2P DBLP:conf/cvpr/SchramowskiBDK23 dataset. Detection of inappropriate content is done by Q16 schramowski2022can.
@@
Method
Locality
CLIP-30K($\uparrow$)
FID-30K($\downarrow$)
@@
Table 15: Quantiative evaluation of concrete object erasure. The best results are highlighted in bold, second-best are underlined. Results of other methods are taken from SPM DBLP:conf/cvpr/Lyu0HCJ00HD24.
@@
F.3 Qualitative results
@@
We see that our method removes Snoopy well (see Fig. 15 while preserving other concepts well (see Fig.  16, 17, 18, 19, 20). In fact, most of the images of non-related concepts generated with CASteer applied are almost identical to those generated by vanilla SDXL.
@@
Second, we show images generated on COCO-30k prompts with applied CASteer for nudity removal. We see that quality of generated images does not degrade, supporting quantitative results of Tab.14. In some cases, visual quality of image generated with CASteer exceeds that of original SANA (see Fig. 22, 3rd and 5th images for example).
@@
Figure 15: Images generated with “Snoopy" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
@@
Figure 16: Images generated with “Mickey" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
@@
Figure 17: Images generated with “Spongebob" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
@@
Figure 18: Images generated with “Pikachu" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
@@
Figure 19: Images generated with “dog" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
@@
Figure 20: Images generated with “legislator" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
@@
Figure 21: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity".
@@
Figure 22: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity".
@@
Figure 23: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity".
Figure 24: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity".
Appendix G Ablations on hyperparameters
@@
In this section, we provide ablations on such hyperparameters as number of prompt pairs to form a steering vector, a value of $\beta$ and choice of the intermediate layer to steer.
G.1 Number of steering vectors
@@
First, we compute steering vectors for the concept of “Snoopy" using SD-1.4 model on varying number of prompt pairs. Then we fix $\beta=2$ and apply CASteer on SD-1.4 model using computed steering vectors on prompts containing “Snoopy" and “Mickey" concepts as described in sec. 4. In particular, we augment each concept using $80$ CLIP DBLP:conf/icml/RadfordKHRGASAM21 templates, and generate $10$ for each concept-template pair, so that for each concept there are $800$ images. We calculate CLIP Score (CS) DBLP:conf/emnlp/HesselHFBC21 and FID DBLP:conf/nips/HeuselRUNH17 on these generated images as described in sec. 4. Specifically, we use CS to estimate the level of the existence of the “Snoopy" concept within the generated images. Next, we calculate FID DBLP:conf/nips/HeuselRUNH17 scores between the set of original generations of SD-1.4 model on “Mickey" prompts and a set of generations of the steered model on “Mickey" prompts. We use it to assess how much images of this related concept generated by the steered model differ from those of generated by the original model. Higher FID value demonstrate more severe generation alteration.
@@
For each number of prompt pairs, we compute the steering vector three times using different non-intersecting sets of prompts. Thus, for each number of prompt pairs, we report three metric values.
@@
Figures 25(a) and 25(b) show CS and FID metrcis for different numbers of prompt pairs. We see that using number of pairs 50 and above results in similar performance.
@@
Figure 25: Ablation on number of prompts for computing steering vectors on SD-1.4. (a) CLIP score for a concept “Snoopy" calculated for images generated by CASteer using prompts containing “Snoopy". (b) FID between original model generations for a concept “Mickey", and generations of a steered model. Blue line indicates mean values across three samples
G.2 Steering strength
@@
Note that CASteer has a hyperparameter $\beta$, which determines the strength of steering for concept removal. Varying values of $\beta$ can lead to different trade-offs between level of target concept erasure and alteration of generated images not containing target concept. Although we observe that value of $\beta=2$ is optimal for all erasing scenarios and we use $\beta=2$ in all our experiments, $\beta$ still can be tuned for each use case. To show how $\beta$ influences performance, we provide results on “Snoopy" erasure using CASTeer on SD-1.4 with different values of $\beta$
We use CASteer on SD-1.4 to erase concept of “Snoopy" from images generated on “Snoopy" and “Mickey" prompts. We calculate CLIP Score (CS) and FID as described above in sec.G.1. Fig. 26(a),26(b) shows the results. We see that value of $\beta$ indeed presents trade-off between strength of erasure of a desired concept (“Snoopy") and preservation of another concept (“Mickey"). Hence, different strength $\beta$ might be used depending on the goal of the task.
Figure 26: Ablation on strength value $\beta$ for steering on SD-1.4 model for erasure. (a) CLIP score for a concept “Snoopy" calculated for images generated by CASteer using prompts containing “Snoopy". (b) FID between original model generations for a concept “Mickey", and generations of a steered model.
G.3 Steering other layers
@@
We use CASteer on SD-1.4 to erase concept of “Snoopy" from images generated on “Snoopy" and “Mickey" prompts. We calculate CLIP Score (CS) and FID as described above in sec.G.1. Tab.16 presents the results. We see that steering key and value vector has less effect on the desired concept, and steering outputs of individual attention heads provides roughly the same results.
@@
G.4 Steering only fraction of layers
@@
We evaluate CASteer on SDXL with $\beta=2$ under these settings for erasing the concept of angriness(fig. 27,  28,  29). In all the figures we illustrate results for values of $k$ between 0 and 60 with a step of 3 for compactness.
Our empirical findings suggest the following: 1) It is not sufficient to steer only one layer (see fig.  29). It can be seen that he effect of steering any single layer is negligible, not causing the desired effect. 2) There is a trade-off between the level of expression of the desired concept in a resulting image and the alteration of general image layout and features. If we steer most of the layers, the overall layout may change drastically from that of the original image and it may cause in the change of identity or other features in the steered image compared to the original one (see fig. 27,  28). As fig. 28 suggests, steering only few last CA layers of the model results in a good trade-off between removing the unwanted concept and keeping other image details intact.
Figure 27: Results on removing the concept of “angry" using CASteer on the prompt “a realistic colorful portrait of an angry man" with steering only $k$ last CA layers, $60\geqslant k\geqslant 1$. Top left corner: image generated without CASteer, images from top to bottom, left to right: images generated using CASteer with varying $k$
Figure 28: Results on removing the concept of “angry" using CASteer on the prompt “a realistic colorful portrait of an angry man" with steering only $k$ first CA layers, $1\leqslant k\leqslant 60$. Top left corner: image generated without CASteer, images from top to bottom, left to right: images generated using CASteer with varying $k$
Figure 29: Results on removing the concept of “angry" using CASteer on the prompt “a realistic colorful portrait of an angry man" with steering only CA layer number $k$, $1\leqslant k\leqslant 60$. Top left corner: image generated without CASteer, images from top to bottom, left to right: images generated using CASteer with varying $k$
Appendix H Other tasks

### v4

@@ G.1 Experimental setup
G.2 Quantitative results
In this section, we present quantitative results on steering SANA model. We use the same experimental setups as for SD-1.4, described in Sec. 4. Results on SANA model are shown in Tab. 1415161718.
Table 14: Quantitative results on nudity removal based on I2P Schramowski et al. (2023) dataset. Detection of nude body parts is done by Nudenet at a threshold of 0.6. F: Female, M: Male. The best results are highlighted in bold, second-best are underlined.
@@ G.2 Quantitative results
Table 15: Quantitative results on inappropriate content removal based on I2P Schramowski et al. (2023) dataset. Detection of inappropriate content is done by Q16 Schramowski et al. (2022).
@@
Method Locality
CLIP-30K($\uparrow$) FID-30K($\downarrow$)
@@
Table 17: Quantiative evaluation of concrete object erasure.
@@
Table 18: Comparison of Artist Concept Removal tasks on SANA model: Famous (left) and Modern artists (right).
Remove “Van Gogh" Remove “Kelly McKernan"
Method $_{e}\uparrow LPIPS$$_{u}\downarrow$ $_{e}\downarrow Acc$$_{u}\uparrow$ $_{e}\uparrow LPIPS$$_{u}\downarrow$ $_{e}\downarrow Acc$$_{u}\uparrow$
SANA - - 1.00 0.89 - - 0.70 0.625
Ours (w/o clip) 0.42 0.14 0.25 0.88 0.19 0.11 0.40 0.620
Ours (clip) 0.42 0.14 0.25 0.875 0.19 0.08 0.30 0.620
G.3 Qualitative results
@@
We see that our method removes Snoopy well (see Fig. 18 while preserving other concepts well (see Fig.  19, 20, 21, 22, 23). In fact, most of the images of non-related concepts generated with CASteer applied are almost identical to those generated by vanilla SANA.
@@
Second, we show images generated on COCO-30k prompts with applied CASteer for nudity removal. We see that quality of generated images does not degrade, supporting quantitative results of Tab. 16 (see Fig. 242526). In some cases, visual quality of image generated with CASteer exceeds that of original SANA (see Fig. 25, 3rd and 5th images for example).
Next, on Fig. 27 and Fig. 28 we show examples of CASteer applied for style removal.
@@
Figure 18: Images generated with “Snoopy” prompts with different seeds. Top: original SANA, bottom: CASteer applied for removing the concept of “Snoopy”.
@@
Figure 19: Images generated with “Mickey” prompts with different seeds. Top: original SANA, bottom: CASteer applied for removing the concept of “Snoopy”.
@@
Figure 20: Images generated with “Spongebob” prompts with different seeds. Top: original SANA, bottom: CASteer applied for removing the concept of “Snoopy”.
@@
Figure 21: Images generated with “Pikachu” prompts with different seeds. Top: original SANA, bottom: CASteer applied for removing the concept of “Snoopy”.
@@
Figure 22: Images generated with “dog” prompts with different seeds. Top: original SANA, bottom: CASteer applied for removing the concept of “Snoopy”.
@@
Figure 23: Images generated with “legislator” prompts with different seeds. Top: original SANA, bottom: CASteer applied for removing the concept of “Snoopy”.
@@
Figure 24: Images generated with different COCO-30k prompts with different seeds. Top: original SANA, bottom: CASteer applied for removing the concept of “nudity”.
@@
Figure 25: Images generated with different COCO-30k prompts with different seeds. Top: original SANA, bottom: CASteer applied for removing the concept of “nudity”.
@@
Figure 26: Images generated with different COCO-30k prompts with different seeds. Top: original SANA, bottom: CASteer applied for removing the concept of “nudity”.
Figure 27: Examples of style removal for SANA. Top: original SANA, bottom: CASteer applied for removing the style of “Van Gogh
Figure 28: Examples of style removal for SANA. Top: original SANA, bottom: CASteer applied for removing the style of “mosaic style” (left 4 images) and “origami style” (right 4 images)
Appendix H Ablations on hyperparameters
@@
In this section, we provide ablations on such hyperparameters as number and choice of prompt pairs to form a steering vector, a value of $\beta$, choice of the intermediate layer to steer, and importance of dot product weighting.
H.1 Number and choice of steering vectors
@@
First, we compute steering vectors for the concept of “Snoopy" using SD-1.4 model on varying number of prompt pairs. Then we fix $\beta=2$ and apply CASteer on SD-1.4 model using computed steering vectors on 800 prompts as described in Sec. 4. In particular, we augment each concept using $80$ CLIP Radford et al. (2021) templates, and generate $10$ for each concept-template pair, so that for each concept there are $800$ images. We calculate CLIP Score (CS) Hessel et al. (2021) and FID Heusel et al. (2017) on these generated images as described in Sec. 4. Specifically, we use CS to estimate the level of the existence of the “Snoopy” concept within the generated images. Next, we calculate average FID Heusel et al. (2017) scores between the set of original generations of SD-1.4 model on prompts non-related to “Snoopy” and a set of generations of the steered model on these prompts. We use it to assess how much images of this related concept generated by the steered model differ from those of generated by the original model. Higher FID value demonstrate more severe generation alteration.
@@
For each number of prompt pairs, we compute the steering vector three times using different non-intersecting sets of prompts and different random seeds. Thus, for each number of prompt pairs, we report three metric values.
@@
Figures 29(a) and 29(b) show CS and FID metrcis for different numbers of prompt pairs used for generation of steering vectors. We see that CASteer performance remains stable across different prompt sets, and using number of pairs 50 and above results in similar performance.
@@
Figure 29: Ablation on number of prompts for computing steering vectors on SD-1.4. (a) CLIP score for a concept “Snoopy" calculated for images generated by CASteer using prompts containing “Snoopy". (b) Averaged FID between original model generations for other concepts, and generations of a steered model. Blue line indicates mean values across three samples
H.2 Importance of dot product weighting
Our choice of steering vector weighting for concept erasure in Eq. 4 is motivated by an observation that dot product between Cross-Attention (CA) output $c$ tokens and the steering vector $s^{X}$ for a concept $X$ serves as a good measure of amount of $X$ that is present in each token of $c$. In this section, we provide experimental ablation on importance of his weighting. We fix $\alpha$ in Eq. 3 to be constant, and perform erasure of “Snoopy" concept using setup as described in Sec. 4. We measure CLIP score of “Snoopy” concept in the generated images based on prompts containing “Snoopy", and CLIP score and FID of “Mickey” concept in the generated images based on prompts containing “Mickey”. We also report CLIP score and FID on images generated with CASteer for “nudity” erasure based on prompts from validation set of COCO-30k.
Tab. 1920 show results for $\alpha=1,2$ with and without clipping applied. We also report results of CASteer as reference. We see that for both values of $\alpha$, “Snoopy" prompt is not erased well. Meanwhile, for $\alpha=2$, image fidelity and prompt alignment suffer, with FID being 3.5 times higher than that of CASteer. This suggests that our proposed weighting mechanism is crusial for performance of our method.
Table 19: Quantiative evaluation of concrete object erasure. The best results are highlighted in bold, second-best are underlined. Results of other methods are taken from SPMLyu et al. (2024) or reproduced.
Snoopy Mickey
[…]
SD-1.4 78.5 74.7 -
Constant steering $\beta=1$ (w/o clip) 76.3 74.3 51.5
Constant steering $\beta=1$ (clip) 76.5 74.0 53.4
Constant steering $\beta=2$ (w/o clip) 72.7 70.7 141.1
Constant steering $\beta=2$ (clip) 72.8 70.8 139.5
Dot Product steering (CASteer, w/o clip) 45.8 70.4 93.0
Dot Product steering (CASteer, clip) 48.5 70.4 89.6
Table 20: General quality estimation of images generated by nudity-erased models. CLIP score and FID are calculated on COCO-30k dataset
Method Locality
CLIP-30K($\uparrow$) FID-30K($\downarrow$)
SD-1.4 31.34 14.04
Constant steering $\beta=1$ (w/o clip) 29.56 14.07
Constant steering $\beta=1$ (clip) 30.84 14.05
Constant steering $\beta=2$ (w/o clip) 27.67 55.34
Constant steering $\beta=2$ (clip) 28.22 55.21
Dot Product steering (CASteer, w/o clip) 30.69 13.28
Dot Product steering (CASteer, clip) 30.09 13.02
H.3 Steering strength
@@
Note that CASteer has a hyperparameter $\beta$, which determines the strength of steering for concept removal. While we set $\beta=2$ in all our experiments, and this choice is motivated by the properties of the resulting transformation (see Sec. 4), it is still possible to use CASteer with varying values of strenght $\beta$. Although experimentally we observe that value of $\beta=2$ is optimal for all erasing scenarios and we use $\beta=2$ in all our experiments, $\beta$ still can be tuned for each use case, and varying values of $\beta$ can lead to different trade-offs between level of target concept erasure and alteration of generated images not containing target concept. To show how $\beta$ influences performance, in this section we provide results on “Snoopy” and “nudity” erasure using CASTeer on SD-1.4, SDXL and SANA with different values of $\beta$
Tab. 21 shows image quality metrics obtained using CASteer erasing “nudity” on COCO-30k validation dataset. It can be seen that steering with strength up to 2 does not lower general image quality for all models. Next, for SDXL and SANA models, steering with strength $\beta>2$ up to $\beta=4$ also does not affect general image quality, while for SD-1.4 steering with $\beta>2$ results in degradation of image fidelity.
Tab. 22 and Tab. 23 show results on nudity erasure using varying steering strength $\beta$. We see that $\beta<2$ results in a lesser suppression of the erased concept. In general, we see that as $\beta$ increases, level of concept suppression is increased. However, for SANA with $\beta\geqslant 3$, and for SDXL with $\beta\geqslant 4$, CASteer without clipping starts showing lower rates of erasure. We hypothesise that this happens due to the absence of clipping, when image tokens with negative dot product with steering vector of “nudity" start containing substantial positive amount of “nudity" after steering. This hypothesis is supported by that CASteer with clipping does not show degrading performance with large values of $\beta$.
Table 21: General quality estimation of images generated by CASteer on SD-1.4, SDXL and SANA model with nudity erasure. CLIP score and FID are calculated on COCO-30k dataset
Method SD-1.4 SDXL SANA
CLIP-30K($\uparrow$) FID-30K($\downarrow$) CLIP-30K($\uparrow$) FID-30K($\downarrow$) CLIP-30K($\uparrow$) FID-30K($\downarrow$)
Vanilla model 31.34 14.04 31.53 13.29 29.28 22.65
Ours, $\beta$=1 31.44 13.87 31.56 13.34 29.31 22.71
Ours, $\beta$=1.5 31.22 13.54 31.57 13.21 29.17 22.62
Ours, $\beta$=2 30.69 13.28 31.51 13.56 28.79 22.83
Ours, $\beta$=2.5 30.36 21.06 31.59 13.49 28.80 22.76
Ours, $\beta$=3.0 29.72 28.65 31.60 13.43 28.82 22.88
Ours, $\beta$=4.0 29.21 51.05 31.59 13.49 28.88 22.78
Ours, $\beta$=1 (clip) 31.42 13.69 31.52 13.27 29.34 22.72
Ours, $\beta$=1.5 (clip) 31.36 13.42 31.54 13.23 29.23 22.56
Ours, $\beta$=2 (clip) 31.09 13.02 31.45 13.37 29.01 22.87
Ours, $\beta$=2.5 (clip) 30.74 16.08 31.44 13.35 29.04 22.71
Ours, $\beta$=3.0 (clip) 30.14 20.11 31.44 13.32 20.03 22.87
Ours, $\beta$=4.0 (clip) 29.98 36.54 31.42 13.37 29.02 22.79
Table 22: Quantitative results on nudity removal based on I2P Schramowski et al. (2023) dataset. using CASteer on SANA model with varying strength $\beta$. Detection of nude body parts is done by Nudenet at a threshold of 0.6. F: Female, M: Male. The best results are highlighted in bold, second-best are underlined.
Method Nudity Detection
[…]
SANA 14 0 3 4 0 5 44 21 91
Ours $\beta=1$ 3 0 0 1 1 1 8 6 20
Ours $\beta=1.5$ 3 0 0 2 0 0 3 3 11
Ours $\beta=2$ 3 0 0 1 0 0 3 2 9
Ours $\beta=2.5$ 2 0 0 1 0 0 3 3 9
Ours $\beta=3$ 4 0 1 2 0 0 4 5 16
Ours $\beta=4$ 6 0 0 3 2 0 4 6 21
Ours $\beta=1$ (clip) 1 0 0 0 1 2 8 5 17
Ours $\beta=1.5$ (clip) 0 0 0 1 0 0 2 3 6
Ours $\beta=2$ (clip) 0 0 0 1 0 0 1 0 2
Ours $\beta=2.5$ (clip) 0 0 0 1 0 0 1 0 2
Ours $\beta=3$ (clip) 1 0 0 0 0 0 1 0 2
Ours $\beta=4$ (clip) 1 0 0 0 0 0 0 1 2
Table 23: Quantitative results on nudity removal based on I2P Schramowski et al. (2023) dataset. using CASteer on SDXL model with varying strength $\beta$. Detection of nude body parts is done by Nudenet at a threshold of 0.6. F: Female, M: Male. The best results are highlighted in bold, second-best are underlined.
Method Nudity Detection
[…]
SDXL 85 7 3 2 7 28 84 66 282
Ours $\beta=1$ 12 0 3 2 1 15 25 28 86
Ours $\beta=1.5$ 9 0 0 1 1 9 12 19 51
Ours $\beta=2$ 4 0 0 0 0 6 10 7 27
Ours $\beta=2.5$ 5 0 0 1 0 5 4 5 20
Ours $\beta=3$ 4 0 0 1 0 5 5 1 16
Ours $\beta=4$ 6 0 1 1 0 5 6 1 20
Ours $\beta=1$ (clip) 15 0 1 1 1 16 27 28 89
Ours $\beta=1.5$ (clip) 7 0 1 0 0 10 11 20 49
Ours $\beta=2$ (clip) 5 1 0 1 0 3 9 7 26
Ours $\beta=2.5$ (clip) 2 0 0 1 0 2 4 4 13
Ours $\beta=3$ (clip) 3 0 1 2 0 3 2 2 13
Ours $\beta=4$ (clip) 3 0 0 1 0 1 3 2 10
H.4 Steering other layers
@@
We use CASteer on SD-1.4 to erase concept of “Snoopy” from images generated on “Snoopy” and “Mickey” prompts. We calculate CLIP Score (CS) and FID as described above in Sec. H.1. Tab. 24 presents the results. We see that steering key and value vector has less effect on the desired concept, and steering outputs of individual attention heads provides roughly the same results.
@@
H.5 Steering only fraction of layers
@@
We evaluate CASteer on SDXL with $\beta=2$ under these settings for erasing the concept of angriness (Fig. 30,  31,  32). In all the figures we illustrate results for values of $k$ between 0 and 60 with a step of 3 for compactness.
Our empirical findings suggest the following: 1) It is not sufficient to steer only one layer (see Fig.  32). It can be seen that the effect of steering any single layer is negligible, not causing the desired effect. 2) There is a trade-off between the level of expression of the desired concept in a resulting image and the alteration of general image layout and features. If we steer most of the layers, the overall layout may change drastically from that of the original image and it may cause in the change of identity or other features in the steered image compared to the original one (see Fig. 30,  31). As Fig. 31 suggests, steering only few last CA layers of the model results in a good trade-off between removing the unwanted concept and keeping other image details intact.
Based on our observations, we hypothesize that steering only few last CA layers of the model is sufficient for erasing texture-based concepts (e.g. styles), while not sufficient for erasing concrete concepts (e.g. “Snoopy”). We test this experimentally by applying CASteer to only 36 last layers of SDXL model. We test erasing “Snoopy” concept and style of “Van Gogh”. Results are presented in Tab. 2526. We see that steering only 36 last layers of SDXL model successfully removes style of “Van Gogh”, while changing images of non-target styles less (see $LPIPS_{u}$). For the concept of “Snoopy”, CASteer applied to only the last 36 layers of the model shows significantly worse erasure performance than CASteer applied to all layers.
Table 25: Quantiative evaluation of concrete object erasure.
Snoopy Mickey Spongebob Pikachu Dog Legislator
[…]
SDXL 74.3 73.1 - 75.1 - 72.7 - 66.3 - 60.8
CASteer (last 36 layers) 57.2 70.8 44.0 74.5 35.2 72.7 19.2 66.3 15.0 60.8 17.3
CASteer 48.7 68.5 69.0 73.7 58.4 72.7 27.8 66.4 37.9 60.8 27.4
Table 26: Comparison of "Van Gogh" Removal tasks on SDXL model.
Remove “Van Gogh"
Method $_{e}\uparrow LPIPS$$_{u}\downarrow$ $_{e}\downarrow Acc$$_{u}\uparrow$
SDXL - - 0.95 0.95
Ours (last 36 layers) 0.30 0.12 0.00 0.85
Ours 0.34 0.19 0.00 0.84
Figure 30: Results on removing the concept of “angry” using CASteer on the prompt “a realistic colorful portrait of an angry man” with steering only $k$ last CA layers, $60\geqslant k\geqslant 1$. Top left corner: image generated without CASteer, images from top to bottom, left to right: images generated using CASteer with varying $k$
Figure 31: Results on removing the concept of “angry” using CASteer on the prompt “a realistic colorful portrait of an angry man” with steering only $k$ first CA layers, $1\leqslant k\leqslant 60$. Top left corner: image generated without CASteer, images from top to bottom, left to right: images generated using CASteer with varying $k$
Figure 32: Results on removing the concept of “angry” using CASteer on the prompt “a realistic colorful portrait of an angry man” with steering only CA layer number $k$, $1\leqslant k\leqslant 60$. Top left corner: image generated without CASteer, images from top to bottom, left to right: images generated using CASteer with varying $k$
Appendix I Other tasks